# Equivariant category of wedges

### v4

@@ Abstract.
for the equivariant category of the wedge $X\vee Y$. As a direct application, we have that the wedge $\bigvee_{i=1}^{m}X_{i}$ is $G$-contractible if and only if each $X_{i}$ is $G$-contractible, for each $i=1,\ldots,m$. One further application is to compute the equivariant category of the quotient $X/A$, for a $G$-space $X$ and an invariant subset $A$ such that the inclusion $A\hookrightarrow X$ is $G$-homotopic to a constant map $\overline{x_{0}}:A\to X$, for some $x_{0}\in X^{G}$. Additionally, we also discus the equivariant and invariant topological complexities for wedges. For instance, as applications of our results, we obtain the following equalities:
@@ 1. Introduction
Let $G$ be a compact Hausdorff topological group acting continuously on a Hausdorff space $X$ on the left. The equivariant category of $X$, denoted by $\text{cat}_{G}(X)$, was introduced by Fadell in [9], as a generalization of the classical category, or Lusternik-Schnirelmann category (LS-category), of a space [12]. For a general overview of LS-category, we refer the reader to the survey article by James [11] and the book by Cornea-Lupton-Oprea-Tanré [5]. In [5, Proposition 1.27(2), p. 14] the authors show that for path-connected normal spaces $X,Y$ with non-degenerate basepoints, the classical category of the wedge $X\vee Y$ is,
@@
Similar to the definition of classical category, $\mathrm{cat}_{G}(X)$ is defined as be the least number of open invariant subsets of $X$ that form a covering for $X$, with each open subset being equivariantly contractible to an orbit, rather than a point.
@@
To this end, we extend the approach to LS-category given by Cornea-Lupton-Oprea-Tanré in [5, Section 1.4] to the equivariant setting.
The main results of this work are:
@@
The paper is organized as follows: In Section 2, we begin with a brief review of $G$-spaces, $G$-maps, $G$-homotopy, and $G$-invariant sets. Basic properties of invariant sets are presented in Lemma 2.1. We recall the definition of equivariant category in Definition 2.2. Some examples of equivariant category are shown in Example 2.3. Lemma 2.4 yields necessary conditions for $G$-contractiblity. The $G$-homotopy invariance of the equivariant category is presented in Proposition 2.5. A key property of the equivariant category is presented in Lemma 2.6. We introduce the notion of $y$-connectivity for a $G$-space in Definition 2.7. As shown in Remark 2.8, $G$-contractibility implies $y$-connectivity and, with an additional condition, $G$-connectivity. In Example 2.9, we use the $y$-connectivity property to compute the equivariant category of the standard Hamiltonian action. Proposition 2.10 shows that, under $y$-connectivity, the inclusion map of any $G$-categorical set is $G$-homotopic to a $G$-map with values in the orbit of $y$, and thus the $A$-Lusternik-Schnirelmann $G$-category is a generalization of the equivariant category. Corollary 2.11 shows that, under certain conditions, the nonequivariant category is a lower bound for the equivariant category. We recall the definition of $G$-cofibration and $G$-retract in Definition 2.12. Proposition (1) provides a characterization of $G$-cofibrations in terms of $G$-retracts, and thus we obtain a key property of $G$-cofibrations. We recall the notion of a $G$-well-pointed space in Definition 2.14. Proposition 2.15 shows that any $G$-non-degenerate basepoint admits a “good” open categorical neighborhood. In Definition 2.16, we recall the notion of a $G$-normal space. Proposition 2.17 shows that any normal $G$-space is a $G$-normal space. Proposition (1) presents properties of $G$-normal spaces.
@@ 2. Preliminary results
In this section, we recall some definitions and we fix notations. We follow the standard notation for compact transformations groups as used in [3] and [15]. Let $G$ be a compact Hausdorff topological group acting continuously on a Hausdorff space $X$ on the left. In this case, we say that $X$ is a $G$-space. For each $x\in X$, the isotropy group or stabilizer of $x$, denoted $G_{x}:=\{g\in G:~{}gx=x\}$, is a closed subgroup of $G$, and the set $O(x):=\{gx:~{}g\in G\}$ is called the orbit of $x$, also denoted by $Gx$. There is a homeomorphism from the quotient space $G/G_{x}$ to $O(x)$, which sends $gG_{x}$ to $g^{-1}x$ for each $g\in G$.
The orbit space $X/G$ is the set of equivalence classes determined by the action, endowed with the quotient topology. Since $G$ is compact and $X$ is Hausdorff, the orbit space $X/G$ is also Hausdorff [15, Proposition 3.1-(v), p. 5], and the quotient map $p:X\to X/G$, sending each point to its orbit, is both open and closed [15, Proposition 3.1, p. 22].
If $H$ is a closed subgroup of $G$, then the set $X^{H}:=\{x\in X:~{}hx=x,\text{ for all }h\in H\}$ is called the $H$-fixed point set of $X$. In particular, the set $X^{G}$ is called the fixed point set of $X$. Here, $X^{H}$ is endowed with the subspace topology. Note that $x\in X^{G}$ if and only if $O(x)=\{x\}$, which is equivalent to $G_{x}=G$.
@@
Note that $\text{cat}_{G}(X)\geq\text{cat}(X/G)$ (using the fact that the quotient map $X\to X/G$ is open). The equality holds whenever $X$ is a free metrizable $G$-space (see [4, Proposition 3.5, p. 2304]).
A $G$-space is called $G$-contractible if $\text{cat}_{G}(X)=1$, i.e., the identity map $1_{X}:X\to X$ is $G$-homotopic to a $G$-map $c:X\to X$ such that $c(X)\subset O(x)$, for some $x\in X$.
@@
    Given a $G$-space $X$, the cone $\text{C}(X)$ of $X$, which is the quotient of $X\times[0,1]$ obtained by collapsing $X\times\{1\}$ to one point, is a $G$-space (with the action $g[x,t]:=[gx,t]$). Moreover, $X\times\{1\}\in\left(\text{C}(X)\right)^{G}$. Note that $\text{cat}_{G}\left(\text{C}(X)\right)=1$, for any $G$-space $X$. The $G$-map $H:\text{C}(X)\times[0,1]\to\text{C}(X)$, given by $H\left([x,t],s\right)=[x,(1-s)t+s]$, satisfies $H_{0}=1_{\text{C}(X)}$ and $H_{1}$ is the constant map at the point $X\times\{1\}$.
@@
    Given a $G$-space $X$, the suspension $SX$ of $X$, which is the quotient of $X\times[0,1]$ obtained by collapsing $X\times\{0\}$ to one point and $X\times\{1\}$ to another point, is a $G$-space (with the action $g[x,t]:=[gx,t]$) and $X\times\{0\},X\times\{1\}\in\left(SX\right)^{G}$. Note that $\text{cat}_{G}(SX)\leq 2$ for any $G$-space $X$. Indeed, an open $G$-categorical covering is given by $\{q\left(X\times[0,3/4)\right),q\left(X\times(1/4,1]\right)\}$, where $q:X\times[0,1]\to SX$ is the quotient map.
@@
    Given $x,y\in X$, there are continuous paths $\alpha:[0,1]\to X^{G_{x}}$ and $\beta:[0,1]\to X^{G_{y}}$ such that $\alpha(0)=x$, $\beta(1)=y$, and $G\alpha(1)=G\beta(0)$.
@@
    Let $H$ be a closed subgroup of $G$. Given $x,y\in X^{H}$, there are continuous paths $\alpha,\beta:[0,1]\to X^{H}$ such that $\alpha(0)=x$, $\beta(1)=y$, and $G\alpha(1)=G\beta(0)$.
@@
The following result is well-known and was established in [14]. However, we provide a proof here in our context.
@@
Proof.
Let $\phi:X\to Y$ and $\psi:Y\to X$ be $G$-maps such that $\phi\circ\psi\simeq_{G}1_{Y}$, and let $F:Y\times[0,1]\to Y$ be a $G$-homotopy between $1_{Y}$ and $\phi\circ\psi$.
Let $U\subset X$ be a $G$-categorical set. Define $V:=\psi^{-1}(U)\subset Y$. Since $\psi$ is a $G$-map, $V$ is open and invariant in $Y$. We aim to show that $V$ is $G$-categorical in $Y$.
Denote by $H:U\times[0,1]\to X$ the $G$-homotopy between the inclusion $\text{incl}_{U}:U\hookrightarrow X$ and a $G$-map $c:U\to X$ such that $c(U)\subset O(x)$, for some $x\in X$.
We define a map $T:V\times[0,1]\to Y$ by the piecewise definition:
$T(y,t):=\left\{\begin{array}[]{ll}F(y,2t),&\hbox{$0\leq t\leq 1/2$;}\\ \phi(H(\psi(y),2t-1)),&\hbox{$1/2\leq t\leq 1$.}\end{array}\right.$
This map is well-defined and continuous because, at $t=1/2$, $F(y,1)=(\phi\circ\psi)(y)$, and $\phi(H(\psi(y),0))=\phi(\psi(y))$, $\forall y\in V$. Moreover, $T$ is a $G$-map.
At $t=0$, $T(y,0)=F(y,0)=y,\forall y\in V$, and at $t=1$, $T(y,1)=c^{\prime}(y)$, for all $y\in V$, where $c^{\prime}:=\phi\circ c\circ\psi:V\to Y$. Note that $c^{\prime}$ is a $G$-map and $c^{\prime}(V)\subset O(\phi(x))$. Hence, the map $T$ is a $G$-homotopy between the inclusion $\text{incl}_{V}$ and the $G$-map $c^{\prime}$ with $c^{\prime}(V)\subset O(\phi(x))$. Therefore, $V\subset Y$ is $G$-categorical.
Now, suppose that $\{U_{1},\ldots,U_{k}\}$ is an open $G$-categorical cover of $X$. Then, the sets $\{V_{1},\ldots,V_{k}\}$, defined as $V_{i}:=\psi^{-1}_{i}(U_{i})$, for each $i$, form an open $G$-categorical cover of $Y$. This proves the first part of the statement.
The second part follows immediately. ∎
@@
(cf. [4, Lemma 3.14, p. 2305]) Let $X$ be a $G$-space, and let $x,y\in X$ such that $H:=G_{x}$. There exists a continuous path $\alpha:[0,1]\to X^{H}$ such that $\alpha(0)=x$ and $\alpha(1)\in Gy$ if and only if there exists a $G$-homotopy $F:Gx\times[0,1]\to X$ such that $F_{0}=incl_{Gx}$ and $F_{1}(Gx)\subset Gy$.
@@
From [4, Definition 3.13, p. 2305], a $G$-space $X$ is said to be $G$-connected if the $H$-fixed point set $X^{H}$ is path-connected, for every closed subgroup $H$ of $G$.
@@
Let $G=S^{1}$ act on $X=S^{2}$ by rotation about the $z$-axis. This is known as the standard Hamiltonian action of $S^{1}$ on $S^{2}$ [5, Example 8.39, p. 248]. The fixed point set of $X$ is $X^{G}=\{p_{N},p_{S}\}\neq\emptyset$, where $p_{N}$ is the north pole and $p_{S}$ is the south pole. Hence, $X$ is not $G$-connected. Furthermore, we have $X^{\{e\}}=X$ and $X^{H}=\{p_{N},p_{S}\}$ for any closed subgroup $H\subset G$ such that $H\neq\{e\}$. Therefore, we have the following cases:
$X^{G_{x}}=\begin{cases}\{p_{N},p_{S}\},&\hbox{ if $x\in X^{G}$},\\ S^{2},&\hbox{ if $x\notin X^{G}$}.\end{cases}$
Thus, $X$ is not $y$-connected for any $y\in X$ (and, of course, it is not $G$-contractible). Hence, we conclude that $\text{cat}_{S^{1}}(S^{2})=2$. Its equivariant categorical covering is given by $U_{1}=S^{2}-\{p_{N}\}$ and $U_{2}=S^{2}-\{p_{S}\}$.
Let $X$ be a $G$-space and $A$ be a $G$-invariant set of $X$. From [13, Definition 2.2, p. 103], the $A$-Lusternik-Schnirelmann $G$-category of $X$, denoted by ${}_{A}\text{cat}_{G}(X)$, is the least integer $k$ such that $X$ can be covered by open $G$-invariant sets $U_{1},\ldots,U_{k}$ of $X$ such that each inclusion map $\text{incl}_{U_{i}}:U_{i}\hookrightarrow X$ is $G$-homotopic to a $G$-map $c_{i}:U_{i}\to X$ with $c_{i}(U_{i})\subset A$. If no such integer $k$ exists, we set ${}_{A}\text{cat}_{G}(X)=\infty$. Note that ${}_{O(x)}\text{cat}_{G}(X)\geq\text{cat}_{G}(X)$, for any $x\in X$.
We have the following statement, which shows that under $y$-connectivity, the inclusion map of any $G$-categorical set is $G$-homotopic to a $G$-map with values in the orbit of $y$. As a direct consequence, we have that the $A$-Lusternik-Schnirelmann $G$-category is a generalization of the equivariant category.
@@
Note that Proposition 2.10 generalizes [13, Remark 2.3, p. 103] where the authors considered more restrictive hypotheses.
Additionally, if $X$ is $x_{0}$-connected for some point $x_{0}\in X^{G}$ (equivalently, if $X$ is $G$-connected and $X^{G}\neq\varnothing$), then ${}_{\{x_{0}\}}\text{cat}_{G}(X)\geq\text{cat}(X)$. Hence, Proposition 2.10 implies the following result, which states that the nonequivariant category provides a lower bound for the equivariant category.
Corollary 2.11.
If $X$ is $G$-connected and $X^{G}\neq\varnothing$, then
$\text{cat}(X)\leq\text{cat}_{G}(X).$
@@
    $A$ is a $G$-retract of $X$ if there exists a $G$-map $r:X\to A$ such that $r(a)=a$, for all $a\in A$.
@@
    $\varphi(x)$ ::=:= $\sup_{t\in I}\mid t-proj_{I}r(x,t)\mid,~{}\text{ for }x\in X;$
    $D(x,t)$ ::=:= $proj_{X}r(x,t),~{}\text{ for }x\in X,~{}t\in I.$
@@
    Here, we recall that $G$ acts trivially on $I$ and diagonally on $X\times I$. Therefore, $D$ is well-defined, equivariant, and satisfies the desired properties.
@@
Examples of equivariant cofibrations are usually a CW pair $(X,A)$, where some group $G$ acts on $X$ (cellularly). Thus, non-free $G$-CW-complexes are $G$-well-pointed.
@@
The following statement is the normal version of [4, Lemma 3.12, p. 2305].
@@
Recall that the quotient map $p:X\to X/G$ is closed. If $X$ is normal, then $X/G$ is normal (see [6, Theorem 3.3, p. 145]). Thus, it suffices to show that the normality of $X/G$ implies the $G$-normality of $X$. Let $A,B\subset X$ be disjoint closed invariant sets. Using the quotient map $p:X\to X/G$, we see that $p(A),p(B)\subset X/G$ are disjoint closed sets because $A$ and $B$ are disjoint invariant sets. Since $X/G$ is normal, there exist disjoint open sets $U,V\subset X/G$ such that $p(A)\subset U$ and $p(B)\subset V$.
@@
Note that any $G$-normal space satisfies the following properties, which are an extension of the nonequivariant case (cf. [6, Theorem 6.1, pg. 152]).
@@
    Recall that $X-U$ is invariant because $U$ is invariant. Since $X$ is $G$-normal, and $A$ and $B:=X-U$ are disjoint closed invariant subsets of $X$ (with $U$ being nonempty), there exist disjoint open invariant sets $F$ and $E$ such that $A\subset F$ and $B\subset E$. Define $V:=F$. Then, we have $\overline{V}\subset U$, as required.
@@
    ()($\Longrightarrow$)( ⟹ ): For each $x\in X$, define $h(x):=\max\{i:~{}x\in U_{i}\}$. We will define the sets $\{V_{i}\}_{i=1}^{n}$ by induction on $n$.

### v5

# Equivariant category and Topological complexity of wedges

@@ Abstract.
for the equivariant category of the wedge $X\vee Y$. As a direct application, we have that the wedge $\bigvee_{i=1}^{m}X_{i}$ is $G$-contractible if and only if each $X_{i}$ is $G$-contractible, for each $i=1,\ldots,m$. One further application is to compute the equivariant category of the quotient $X/A$, for a $G$-space $X$ and an invariant subset $A$ such that the inclusion $A\hookrightarrow X$ is $G$-homotopic to a constant map $\overline{x_{0}}:A\to X$, for some $x_{0}\in X^{G}$. Additionally, we discuss the equivariant and invariant topological complexities for wedges. For instance, as applications of our results, we obtain the following equalities:
@@ 1. Introduction
Let $G$ be a compact Hausdorff topological group acting continuously on a Hausdorff space $X$ on the left. The equivariant category of $X$, denoted by $\text{cat}_{G}(X)$, was introduced by Fadell in [10], as a generalization of the classical category, or Lusternik-Schnirelmann category (LS-category), of a space [13]. For a general overview of LS-category, we refer the reader to the survey article by James [12] and the book by Cornea-Lupton-Oprea-Tanré [6]. In [6, Proposition 1.27(2), p. 14] the authors show that for path-connected normal spaces $X,Y$ with non-degenerate basepoints, the classical category of the wedge $X\vee Y$ is,
@@
Similar to the definition of classical category, $\mathrm{cat}_{G}(X)$ is defined as the least number of open invariant subsets of $X$ that form a covering for $X$, with each open subset being equivariantly contractible to an orbit, rather than a point.
@@
To this end, we extend the approach to LS-category given by Cornea-Lupton-Oprea-Tanré in [6, Section 1.4] to the equivariant setting.
Additionally, we discuss the equivariant and invariant topological complexities for wedge spaces.
In simpler terms, the equivariant topological complexity $\mathrm{TC}_{G}$ is the minimal number of $G$-equivariant motion planning rules, whereas the invariant topological complexity $\mathrm{TC}^{G}$ is the minimal number of $G\times G$-equivariant motion planning rules required to navigate the space in a way that respects the group symmetry (see Definition 4.1).
The main results of this work are the following.
@@
The paper is organized as follows: In Section 2, we begin with a brief review of $G$-spaces, $G$-maps, $G$-homotopy, and $G$-invariant sets. Basic properties of invariant sets are presented in Lemma 2.1. We recall the definition of equivariant category in Definition 2.2. Some computation of equivariant category are shown in Example 2.3. Lemma 2.4 yields necessary conditions for $G$-contractiblity. The $G$-homotopy invariance of the equivariant category is presented in Proposition 2.5. A key property of the equivariant category is presented in Lemma 2.6. We introduce the notion of $y$-connectivity for a $G$-space in Definition 2.7. As shown in Remark 2.8, $G$-contractibility implies $y$-connectivity and, with an additional condition, $G$-connectivity. In Example 2.9, we use the $y$-connectivity property to compute the equivariant category of the standard Hamiltonian action. Proposition 2.10 shows that, under $y$-connectivity, the inclusion map of any $G$-categorical set is $G$-homotopic to a $G$-map with values in the orbit of $y$. We recall the definition of $G$-cofibration and $G$-retract in Definition 2.11. Proposition (1) provides a characterization of $G$-cofibrations in terms of $G$-retracts, and thus we obtain a key property of $G$-cofibrations. We recall the notion of a $G$-well-pointed space in Definition 2.13. Proposition 2.14 shows that any $G$-non-degenerate basepoint admits a “good” open categorical neighborhood. In Definition 2.15, we recall the notion of a $G$-normal space. Proposition 2.16 shows that any normal $G$-space is a $G$-normal space. Proposition (1) presents properties of $G$-normal spaces.
@@
The authors of this paper deeply thank the referees for very valuable comments and timely corrections on previous versions of the work.
@@ 2. Preliminary results
In this section, we recall some definitions and fix our notation. We follow the standard conventions for compact transformation groups as used in [4] and [17].
Let $G$ be a compact Hausdorff topological group acting continuously on a Hausdorff space $X$ on the left. In this case, we say that $X$ is a $G$-space.
For each $x\in X$, the isotropy group (or stabilizer) of $x$, denoted
$G_{x}:=\{g\in G:gx=x\},$
is a closed subgroup of $G$.
The set
${O}(x):=\{gx:g\in G\},$
also denoted by $Gx$, is called the orbit of $x$.
There exists a homeomorphism from the quotient space $G/G_{x}$ to O(x){O}(x), which maps $gG_{x}$ to $g^{-1}x$, for each $g\in G$.
The orbit space $X/G$ is the set of equivalence classes determined by the action, endowed with the quotient topology. Since $G$ is compact and $X$ is Hausdorff, the orbit space $X/G$ is also Hausdorff [17, Proposition 3.1-(v), p. 5], and the quotient map $p:X\to X/G$, sending each point to its orbit, is both open and closed [17, Proposition 3.1, p. 22].
If $H$ is a closed subgroup of $G$, then the set
$X^{H}:=\{x\in X:~hx=x,\text{ for all }h\in H\}$
is called the $H$-fixed point set of $X$. In particular, the set $X^{G}$ is called the fixed point set of $X$. Here, $X^{H}$ is endowed with the subspace topology. Note that $x\in X^{G}$ if and only if $O(x)=\{x\}$, which is equivalent to $G_{x}=G$.
@@
Note that $\text{cat}_{G}(X)\geq\text{cat}(X/G)$ (using the fact that the quotient map $X\to X/G$ is open). The equality holds whenever $X$ is a free metrizable $G$-space (see [5, Proposition 3.5, p. 2304]).
A $G$-space $X$ is called $G$-contractible if $\text{cat}_{G}(X)=1$, i.e., the identity map $1_{X}:X\to X$ is $G$-homotopic to a $G$-map $c:X\to X$ such that $c(X)\subset O(x)$, for some $x\in X$.
@@
    Given a $G$-space $X$, the cone $C(X)$ of $X$, defined as the quotient of $X\times[0,1]$ obtained by collapsing $X\times\{1\}$ to a single point, is a $G$-space, with the action defined by $g[x,t]:=[gx,t]$. Moreover, the point $X\times\{1\}$ belongs to the fixed point set $(C(X))^{G}$. Note that $\operatorname{cat}_{G}(C(X))=1$ for any $G$-space $X$. The $G$-map $H:C(X)\times[0,1]\to C(X)$, given by
    H([x,t],s)=[x,(1s)t+s],H([x,t],s)=[x,(1-s)t+s],
    satisfies $H_{0}=\operatorname{id}_{C(X)}$ and $H_{1}$ is the constant map at the point $X\times\{1\}$.
@@
    Given a $G$-space $X$, the suspension $SX$ of $X$, defined as the quotient of $X\times[0,1]$ obtained by collapsing $X\times\{0\}$ to one point and $X\times\{1\}$ to another, is itself a $G$-space, with the action given by $g[x,t]:=[gx,t]$. Moreover, the points $X\times\{0\}$ and $X\times\{1\}$ belong to the fixed point set $(SX)^{G}$. Note that $\operatorname{cat}_{G}(SX)\leq 2$ for any $G$-space $X$. Indeed, an open $G$-categorical covering is given by
    $\left\{q\left(X\times[0,3/4)\right),\;q\left(X\times(1/4,1]\right)\right\},$
    where $q:X\times[0,1]\to SX$ is the quotient map.
@@
    Given $x,y\in X$, there exist continuous paths $\alpha:[0,1]\to X^{G_{x}}$ and $\beta:[0,1]\to X^{G_{y}}$ such that $\alpha(0)=x$, $\beta(1)=y$, and $G\alpha(1)=G\beta(0)$.
@@
    Let $H$ be a closed subgroup of $G$. Given $x,y\in X^{H}$, there exist continuous paths $\alpha,\beta:[0,1]\to X^{H}$ such that $\alpha(0)=x$, $\beta(1)=y$, and $G\alpha(1)=G\beta(0)$.
@@
The following result is well-known and was established in [15].
@@
(cf. [5, Lemma 3.14, p. 2305]) Let $X$ be a $G$-space, and let $x,y\in X$ such that $H:=G_{x}$. There exists a continuous path $\alpha:[0,1]\to X^{H}$ such that $\alpha(0)=x$ and $\alpha(1)\in Gy$ if and only if there exists a $G$-homotopy $F:Gx\times[0,1]\to X$ such that $F_{0}=incl_{Gx}$ and $F_{1}(Gx)\subset Gy$.
@@
From [5, Definition 3.13, p. 2305], a $G$-space $X$ is said to be $G$-connected if the $H$-fixed point set $X^{H}$ is path-connected, for every closed subgroup $H$ of $G$.
@@
Let $G=S^{1}$ act on $X=S^{2}$ by rotation about the $z$-axis. This is known as the standard Hamiltonian action of $S^{1}$ on $S^{2}$ [6, Example 8.39, p. 248]. The fixed point set of $X$ is $X^{G}=\{p_{N},p_{S}\}\neq\emptyset$, where $p_{N}$ is the north pole and $p_{S}$ is the south pole. Hence, $X$ is not $G$-connected. Furthermore, $X$ is not $y$-connected for any $y\in X$ (and, of course, it is not $G$-contractible). Hence, we conclude that $\text{cat}_{S^{1}}(S^{2})=2$. Its equivariant categorical covering is given by $U_{1}=S^{2}-\{p_{N}\}$ and $U_{2}=S^{2}-\{p_{S}\}$.
Let $X$ be a $G$-space and $A$ be a $G$-invariant set of $X$. From [14, Definition 2.2, p. 103], the $A$-Lusternik-Schnirelmann $G$-category of $X$, denoted by $_{A}\text{cat}_{G}(X)$, is the least integer $k$ such that $X$ can be covered by open $G$-invariant sets $U_{1},\ldots,U_{k}$ of $X$ such that each inclusion map $\text{incl}_{U_{i}}:U_{i}\hookrightarrow X$ is $G$-homotopic to a $G$-map $c_{i}:U_{i}\to X$ with $c_{i}(U_{i})\subset A$. If no such integer $k$ exists, we set $_{A}\text{cat}_{G}(X)=\infty$. Note that $_{O(x)}\text{cat}_{G}(X)\geq\text{cat}_{G}(X)$, for any $x\in X$.
Further discussion of $_{A}\text{cat}_{G}(X)$ can be found in [14] and [3].
We have the following statement, which shows that under $y$-connectivity, the inclusion map of any $G$-categorical set is $G$-homotopic to a $G$-map with values in the orbit of $y$. As a direct consequence, the $A$-Lusternik-Schnirelmann $G$-category satisfies the following property: there exists a subset $A\subseteq X$ such that $_{A}cat_{G}(X)=cat_{G}(X)$, whenever $X$ is $y$-connected.
@@
Note that Proposition 2.10 generalizes [14, Remark 2.3, p. 103], where the authors considered more restrictive hypotheses.
Additionally, if $X$ is $x_{0}$-connected for some point $x_{0}\in X^{G}$ (equivalently, if $X$ is $G$-connected and $X^{G}\neq\varnothing$), then $_{\{x_{0}\}}\text{cat}_{G}(X)\geq\text{cat}(X)$. Hence, Proposition 2.10 implies that the nonequivariant category provides a lower bound for the equivariant category, i.e., $\text{cat}(X)\leq\text{cat}_{G}(X)$ for any $G$-connected space $X$ with $X^{G}\neq\varnothing$.
@@
    AA is a $G$-retract of $X$ if there exists a $G$-map $r:X\to A$ such that $r(a)=a$, for all $a\in A$. Such a map $r$ is called a $G$-retraction.
@@
    $\varphi(x)$ $:=$ $\sup_{t\in I}\mid t-proj_{I}r(x,t)\mid,~\text{ for }x\in X;$
    $D(x,t)$ $:=$ $proj_{X}r(x,t),~\text{ for }x\in X,~t\in I.$
@@
    Here, we recall that $G$ acts trivially on $I$ and diagonally on $X\times I$. Furthermore, $D$ is well-defined, equivariant, and satisfies the desired properties. Note that for $t>\varphi(x)$, we have $r(x,t)\in A\times I$, since $t>0$. Thus, $D(x,t)\in A$.
@@
Examples of equivariant cofibrations are usually a CW pair $(X,A)$, where some group $G$ acts on $X$ (cellularly) [16, Theorem 3.1]. Thus, every $G$-CW-complex with a $G$-fixed basepoint is $G$-wellpointed, since the inclusion of the basepoint is a $G$-cofibration.
@@
The following statement is the normal version of [5, Lemma 3.12, p. 2305].
@@
Recall that the quotient map $p:X\to X/G$ is closed. If $X$ is normal, then $X/G$ is normal (see [7, Theorem 3.3, p. 145]). Thus, it suffices to show that the normality of $X/G$ implies the $G$-normality of $X$. Let $A,B\subset X$ be disjoint closed invariant sets. Using the quotient map $p:X\to X/G$, we see that $p(A),p(B)\subset X/G$ are disjoint closed sets because $A$ and $B$ are disjoint invariant sets. Since $X/G$ is normal, there exist disjoint open sets $U,V\subset X/G$ such that $p(A)\subset U$ and $p(B)\subset V$.
@@
Note that any $G$-normal space satisfies the following properties, which are an extension of the nonequivariant case (cf. [7, Theorem 6.1, pg. 152]).
@@
    Recall that XUX-U is invariant because $U$ is invariant. Since $X$ is $G$-normal, and $A$ and B:=XUB:=X-U are disjoint closed invariant subsets of $X$ (with $U$ being nonempty), there exist disjoint open invariant sets $F$ and $E$ such that $A\subset F$ and $B\subset E$. Define $V:=F$. Then, we have $\overline{V}\subset U$, as required.
@@
    $(\Longrightarrow)$: For each $x\in X$, define $h(x):=\max\{i:~x\in U_{i}\}$. We will define the sets $\{V_{i}\}_{i=1}^{n}$ by induction on $n$.